Let C(x)=(1+x^2)^3. Find the functions f and g so that C(x)=(f o g)(x) is true for all values of x.
Answer by stanbon(75887)

Let C(x)=(1+x^2)^3. Find the functions f and g so that C(x)=(f o g)(x) is true for all values of x.
------------------
g(x) = 1+x^2
f(x) = x^3
-----
fog(x) = f(g(x)) = f(1+x^2) = (1+x^2)^3
----------------------
Cheers,
Stan H.
------------